### v3

@@ Abstract
Reference models convey best practices and standards. The reference frameworks necessitate conformance checks to ensure adherence to established guidelines and principles, which is crucial for maintaining quality and consistency in various processes. This paper explores automated conformance checks for concrete process models against reference models using causal dependency analysis of tasks and events. Existing notions of conformance checking for process models focus on verifying process execution traces and lack the expressiveness and automation needed for semantic model comparison, leaving this question unresolved. We integrate our approach into a broader semantic framework for defining reference model conformance. We outline an algorithm for reference process model conformance checking, evaluate it through a case study, and discuss its strengths and limitations. Our research provides a tool-assisted solution enhancing accuracy and flexibility in process model conformance verification.
@@ 3 Background and Related Work
A reference model is defined as an abstract representation within a given modeling language that captures domain concepts and relationships, specifying properties that must hold for any conformant concrete model, e.g., which model elements may or must exist and how they interrelate [23]. A robust conformance relation must guarantee that all relevant semantic properties are preserved by conformant concrete models Konersmann et al. [13]. More specifically, a conformant concrete model must contain appropriate incarnations of relevant elements from the reference model. These incarnations are concrete model elements that correspond to reference elements while preserving their interrelations in the concrete model.
@@
To enable concrete models to extend a reference model in a meaningful manner, it is generally more appropriate to adopt an open-world assumption regarding model semantics for conformance relations, rather than a closed-world assumption. This perspective allows a concrete model to incorporate additional elements that do not have counterparts in the reference model while still maintaining conformance. In terms of process model semantics, this implies that the execution trace of the concrete model may include additional tasks and events, provided that the causal relationships among tasks and events in the reference model are respected in their respective implementations. It is essential to note that we assume tasks and events are uniquely identified by names within a process model, ensuring that each task and event appears only once.
@@ 4 Conformance Checking Approach
We define the open-world semantics of a process model as the collection of process execution traces that uphold the local causal dependencies of tasks and events as specified by the process model. In this context, each instance of a task or event is permitted to occur only after a suitable configuration of its predecessor instances has taken place and before an appropriate configuration of its successor instances is realized. Furthermore, these configurations of predecessor and successor instances must also be maintained among instances of the same task or event. In other words, loops that disrupt the local dependency structure of tasks and events are explicitly disallowed.
@@
This approach bears similarity to our conformance checking methodology for class diagrams as described in [13], where an incarnation is deemed conformant to its corresponding reference element if its properties and relationships with other elements are preserved. However, in the context of concrete process models, it is inadequate to limit our examination to neighboring tasks and events. For instance, consider the following scenarios:
@@
  • Violation of Exclusive Alternatives: The relationships defined by exclusive alternatives among reference tasks may be contravened in the concrete model if their incarnations are executed sequentially. In such cases, the concrete model fails to be a refinement of the reference model and thus does not conform to it.
@@
The conformance checking algorithm is divided into two distinct phases. In the first phase, we identify the local causal dependencies of tasks and events within the reference model and encode these dependencies into propositional logic formulas—one for all direct predecessors and another for all direct successors. In the second phase, we perform the conformance check on the concrete model by analyzing the causal relationships of each incarnation of a task or event with their predecessors and successors in the concrete model, subsequently comparing these relationships to the local causal dependencies expressed in the corresponding formulas.

### v4

@@ Abstract
Reference models convey best practices and standards crucial for maintaining quality and consistency in various processes. Conformance checks are necessary to ensure adherence to these models and the established guidelines and principles they encode. This paper explores automated conformance checking for concrete process models against reference models using causal dependency analysis of tasks and events. Existing notions of conformance checking for process models focus on verifying process execution traces and lack the expressiveness and automation needed for semantic model comparison, leaving this question unresolved. Our approach is integrated into a broader semantic framework for defining reference model conformance. We outline an algorithm for reference process model conformance checking, evaluate its implementation through a case study, and discuss its strengths and limitations. Our research provides a tool-assisted solution enhancing accuracy and flexibility in process model conformance verification.
@@ 3 Background and Related Work
A reference model is defined as an abstract representation within a given modeling language that captures domain concepts and their domain-specific relations, specifying properties that must hold for any conformant concrete model, e.g., which model elements may or must exist and how they interrelate [23]. A robust conformance relation must guarantee that all relevant semantic properties are preserved by conformant concrete models Konersmann et al. [13]. More specifically, a conformant concrete model must contain appropriate incarnations of relevant elements from the reference model. These incarnations are concrete model elements that correspond to reference elements while preserving their interrelations in the concrete model.
@@
To enable concrete models to extend a reference model in a meaningful manner, it is generally more appropriate to adopt an open-world assumption regarding model semantics for conformance relations, rather than a closed-world assumption. This perspective allows a concrete model to incorporate additional elements that do not have counterparts in the reference model while still maintaining conformance. In terms of process model semantics, this implies that the execution trace of the concrete model may include additional tasks and events, provided that the causal relations among tasks and events in the reference model are preserved by their respective incarnations. It is essential to note that we assume tasks and events are uniquely identified by names within a process model, ensuring that each task and event appears only once.
@@ 4 Conformance Checking Approach
We define the open-world semantics of a process model as the collection of process execution traces that uphold the local causal dependencies of tasks and events as specified by the process model. In this context, each instance of a task or event is permitted to occur only after a suitable configuration of its predecessor instances has taken place and before an appropriate configuration of its successor instances is realized. Furthermore, these configurations of predecessor and successor instances must also be maintained between instances of the same task or event. In other words, loops that disrupt the local dependency structure of tasks and events are explicitly disallowed.
@@
This approach bears similarity to our conformance check for class diagrams as described in [13], where an incarnation is deemed conformant to its corresponding reference element if its properties and relationships with other elements are preserved. However, in the context of concrete process models, it is inadequate to limit our examination to neighboring tasks and events. For instance, consider the following scenarios:
@@
  • Violation of Exclusive Alternatives: The causal relations defined by exclusive alternatives among reference tasks may be contravened in the concrete model if their incarnations are executed sequentially. In such cases, the concrete model fails to be a refinement of the reference model and thus does not conform to it.
@@
The conformance checking algorithm is divided into two distinct phases. In the first phase, we identify the local causal dependencies of tasks and events within the reference model and encode these dependencies into propositional logic formulas—one for all direct predecessors and another for all direct successors. In the second phase, we perform the conformance check on the concrete model by analyzing the causal relation of each incarnation of a task or event with their predecessors and successors in the concrete model, subsequently comparing these relations to the local causal dependencies expressed in the corresponding formulas.